### v1

# Equivariant concordance of periodic 2-knots in $S^{4}$

Abstract.
We show that the equivariant concordance group of smooth 2-knots in $S^{4}$ invariant under a $\mathbb{Z}/d\mathbb{Z}$ action, where the action is given by rotation about an unknotted sphere intersecting the 2-knot in two points, is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for all $d\geq 2$. This is in contrast to the non-equivariant case, in which all 2-knots are slice. We construct a new invariant for these 2-knots, which we call periodic, and show that it fully classifies them up to equivariant concordance. The invariant depends on an extension of the Arf invariant for null-homologous classical knots in arbitrary 3-manifolds.
Introduction
Concordance of surfaces in various contexts has been an object of interest to topologists as early as the 1960s. In 1965, Kervaire [Ker65] showed that all 2-spheres in $S^{4}$ are slice, meaning that they bound smoothly embedded 3-balls into $B^{5}$. In 2015, Sunukjian [Sun15] generalized this to the statement that the concordance class of any oriented surface in a simply-connected 4-manifold is determined entirely by its genus and homology class. Our work uses many of the same techniques as these earlier papers, and extends the theory of concordance of 2-spheres to a $\mathbb{Z}/d\mathbb{Z}$-equivariant setting.
Our goal is to examine equivariant concordance of smooth 2-spheres that are symmetric with respect to an ambient rotation $\tau$ acting on $S^{4}$ fixing an unknotted $S^{2}$ of fixed finite order $d$, where the rotation restricted to the symmetric 2-sphere fixes exactly two points. In analogy with the classical dimension, we call such 2-knots $d$-periodic. We say that these spheres are equivariantly slice if they bound 3-balls into $B^{5}$ that are themselves equivariant with respect the extension of $\tau$ to $B^{5}$ that fixes an unknotted properly embedded 3-ball. Note that any such slice ball must intersect the fixed set of $\tau$ extended to $B^{5}$ in a single unknotted arc by Morgan’s resolution of the Smith Conjecture [Mor84]. There is also a notion of periodic connected sum, analogous to the connected sum of strongly invertible knots defined by Sakuma [Sak86], which turns the set of smooth concordance classes of periodic 2-knots up to equivariant isotopy into a group, denoted $\mathcal{C}^{2}_{d}$. This paper introduces a new invariant $\alpha$ of $d$-periodic 2-knots which completely determines their equivariant concordance class, simultaneously computing of all of these groups.
Theorem 5.4.
There exists a $\mathbb{Z}/2\mathbb{Z}$-valued, additive concordance invariant $\alpha$ for $d$-periodic 2-knots $S$ such that:
  1. (1)
    if $\alpha(S)=0$, then $S$ is equivariantly slice,
  2. (2)
    if $\alpha(S)=1$, then $S$ is equivariantly concordant to the $d$-twist spun trefoil, and
  3. (3)
    every $d$-periodic 2-knot satisfies exactly one of these conditions.
Thus the smooth equivariant concordance group $\mathcal{C}^{2}_{d}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for all $d$.
We prove our main theorem as follows. We first construct equivariant slice balls by constructing equivariant Seifert 3-manifolds, pushing these into $B^{5}$, and then equivariantly surgering them to 3-balls. The surgery approach fails under certain conditions, which we then show to be due to an actual equivariant concordance obstruction. This obstruction relies on a new extension of the Arf invariant for knots to null-homologous knots in arbitrary 3-manifolds, detailed in Section 3. Realizing this invariant as an equivariant sliceness obstruction requires us to also build equivariant Seifert 4-manifolds for periodic 3-manifolds in ambient 5-space with careful control of the fixed set.
It is worth emphasizing that the concordance group, $\mathbb{Z}/2\mathbb{Z}$, does not depend on the period $d$ of the rotation. In short, this is because the equivariant concordance of periodic spheres depends only on the isotopy class of the quotient sphere relative to its intersection with the axis of rotation. We thus obstruct this type of relative concordance between these quotient spheres, and include criteria for when this concordance can be lifted to an equivariant concordance on the cover. Thus our proof really shows that the concordance group of annuli properly embedded in $S^{1}\times B^{3}$ is $\mathbb{Z}/2\mathbb{Z}$, explaining the lack of dependence on $d$. Another consequence of our proof is that the natural map from $\mathcal{C}^{2}_{kd}$ to $\mathcal{C}^{2}_{d}$ for $k\in\mathbb{N}$ induced by the inclusion of $kd$-periodic 2-knots into the set of $d$-periodic 2-knots is always an isomorphism, which further reflects insignificance of the period.
This result fits into a body of work both on equivariant 2-knots and on surface concordance, in various categories. Ruberman [Rub84] showed that 2-knots which are equivariant with respect to free involutions on $S^{4}$ are unique up to equivariant concordance. Gabbard [Gab24] has also worked on 2-spheres that are also equivariant with respect to a rotation involution on $S^{4}$, but which are embedded in such a way that the rotation induces an orientation-reversing involution on the 2-sphere fixing a circle.
Related also is the work of McCooey [McC07] on topological concordance of $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ actions on $S^{4}$ for primes $p$, where he also found a $\mathbb{Z}/2\mathbb{Z}$-valued action concordance invariant. McCooey’s definition of concordance of actions implies concordance of their fixed sets, and so in proving his theorem, he shows that there are at most two topological concordance classes of pairs of 2-knots in $S^{4}$ intersecting in two points, often called Montesinos twins, each of which are the fixed set of a different commuting $\mathbb{Z}/p\mathbb{Z}$ action on $S^{4}$. His proof uses surgery theory, and his obstruction comes from the generator of the $\mathbb{Z}/2\mathbb{Z}$ part of $L_{4}(\mathbb{Z}\times\mathbb{Z})\cong\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z}$.
Our work is substantially different from that of McCooey, although we conjecture that our obstruction to smooth equivariant sliceness, restricted to the topological category and the correct class of periodic 2-knots, is equivalent to McCooey’s. For one, McCooey’s result only has implications for equivariant topological concordance within the class of 2-knots which are not only $\mathbb{Z}/p\mathbb{Z}$-equivariant, but are also themselves the fixed sets of $\mathbb{Z}/p\mathbb{Z}$ actions on $S^{4}$ for $p$ prime. As mentioned before, his proof is fully surgery-theoretic, and he was unable to find an example of a $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ action on $S^{4}$ for which his obstruction vanished. In contrast, our theorem applies to all smooth periodic 2-knots of any period. The proof is completely constructive, and we give explicit criteria for the vanishing of our obstruction and identify it as a direct extension of the Arf invariant of a knot, as well as examples where it does not vanish. Although the Arf invariant is typically a topological invariant, we work in the smooth category for the remainder of this paper.
There is also a family of literature on surface concordance in non-simply connected 4-manifolds, in which there are several known obstructions. In 2018 Gabai gave a proof of the 4-dimensional lightbulb theorem [Gab20], which gives conditions under which a sphere with a framed, geometrically dual sphere can be unknotted. Soon after, Miller proved a version for sphere concordance in the presence of framed duals [Mil21]. Both of these proofs require that the Freedman-Quinn invariant of the given 2-knot vanish in the case that the ambient 4-manifold has 2-torsion in its fundamental group. The connection to the Freedman-Quinn invariant was formalized by Schneiderman and Teichner [ST22] after a counterexample to the lightbulb theorem in the case of ambient 2-torsion was constructed by Schwartz [Sch19]. In the case that a sphere in a non-simply connected 4-manifold has an unframed dual, Klug and Miller [KM21] extended work of Stong [Sto93] and Freedman and Quinn [FQ90] to show that the Kervaire-Milnor invariant of the sphere can be used to obstruct concordance. Although we do not use these obstructions, we mention them in order to give a sense of scope of surface concordance invariants.
Outline
  • In Section 1, we establish some preliminaries of our setup and outline some necessary results regarding equivariant Morse theory with vertical boundary. This requires the language of Morse theory for manifolds with boundary, as described by Borodzik-Némethi-Ranicki in [BNR16], as well as work of Borodzik-Powell in [BP16] and Schwartz in [Sch80].
  • In Section 2, we prove the existence of equivariant Seifert 3-manifolds for equivariant 2-knots while controlling the fixed set by adapting a technique of Yanagawa in [Yan69].
  • In Section 3, we take an interlude to describe an extension of the Arf invariant to null-homologous knots in 3-manifolds, inspired by a definition of Robertello in [Rob65].
  • In Section 4, we use the extension in Section 3 to define the $\alpha$ invariant for $d$-periodic 2-knots together with their Seifert solids, and then give a procedure to produce equivariant slice balls via surgery on these Seifert solids under the assumption that it vanishes. We also show that all knots for which $\alpha=1$ are also mutually equivariantly concordant.
  • In Section 5, we use the invariants defined in the previous sections to demonstrate that $\alpha$ is an invariant of $d$-periodic 2-knots up to equivariant isotopy, and that it obstructs equivariant sliceness. The proof involves the construction of certain equivariant Seifert 4-manifolds for 3-manifolds with $\mathbb{Z}/d\mathbb{Z}$ actions fixing a circle.
Acknowledgments
The author would like to thank Maggie Miller, Imogen Montague, Mark Powell, Evan Scott, Aru Mukherjea, and Seungwon Kim for their generous help and advice. Special thanks to Malcolm Gabbard for suggesting this question, and to Mark Powell for his helpful comments on an earlier draft of this paper.
1. Preliminaries
1.1. Main definitions and conventions
In this paper, we work entirely in the smooth category. It is important for consistency to fix the following basic convention: when we say that $S$ is a 2-knot in $S^{4}$, we mean that $S$ is the image of a smooth embedding $\psi\colon S^{2}\to S^{4}$, considered up to smooth isotopy of $\psi$.
We begin with some background about smooth group actions on manifolds. We specify to the case of $\mathbb{Z}/d\mathbb{Z}$ for this paper, but it is worth noting that there is a large body of work on 4-manifolds with various group actions; see, for example, Edmonds’ survey [Edm18].
Definition 1.1.
An action $\tau$ of $\mathbb{Z}/d\mathbb{Z}$ on the standard $m$-ball $B^{m}$ is linear if it is conjugate to the action of an $m$-dimensional orthogonal representation of $\mathbb{Z}/dZ$ on $B^{m}$. An action of $\mathbb{Z}/d\mathbb{Z}$ on $S^{m-1}$ is linear if it is conjugate to the action of such a representation on $B^{m}$ restricted to the boundary.
The actions we consider in this paper are the the linear rotation action $\rho$ of $\mathbb{Z}/d\mathbb{Z}$ on $S^{2}$ fixing two points, and the linear rotation action $\tau$ of $\mathbb{Z}/d\mathbb{Z}$ on $B^{5}$ fixing an unknotted 3-ball such that its restriction to the boundary $S^{4}$ fixes an unknotted 2-sphere. Restricting to linear actions simplifies the argument, although we hope to extend to nonlinear actions in future work. For convenience, we give a general definition for embeddings which are equivariant with respect to this type of action in all dimensions.
Definition 1.2.
Let $M$ be a smooth $n$-manifold and let $\psi\colon M\to S^{m}$ be a smooth embedding. We say that $\psi$ is $d$-periodic if there exists an order $d$ linear rotation $\tau$ on $S^{m}$ fixing an unknotted smooth copy of $S^{m-2}\subset S^{m}$ such that $\tau$ fixes the image of $\psi$ setwise and $\psi(M)$ intersects $\operatorname{Fix}(\tau)$ transversely.
In the case that $\psi(M)\cap\operatorname{Fix}(\tau)$ is a single, smoothly embedded copy of $S^{n-2}$, or a single properly embedded copy of $D^{n-2}$ in the case that $M$ has boundary, we say that $\psi$ is simply $d$-periodic.
Some abuses of this notation: where it is not explicitly relevant, we will often drop the prefix $d$- and describe embeddings just as periodic or simply periodic. Furthermore, we will also often describe the image of periodic embeddings as periodic; i.e., when we say that a 2-knot $S\subset S^{4}$ is periodic, we mean that there is an embedding $\psi\colon S^{2}\to S^{4}$ whose image is $S$, and that $\psi$ is periodic for some period $d$.
Following this definition, we see that a 2-knot $S$ in $S^{4}$ is $d$-periodic if it satisfies
$\tau\circ\psi=\psi\circ\rho,$
with $\operatorname{Fix}(\tau)$ an unknotted 2-sphere. In this case, since the restriction of $\tau$ to $S$ is orientation preserving, we see by the Lefschetz fixed point theorem that $S\cap\operatorname{Fix}(\tau)$ consists of exactly two points. Such spheres arise naturally in common constructions of 2-knots: all $d$-twist spun knots can be constructed to have this symmetry about their spin axis.
Remark 1.3.
All periodic 2-knots are simply periodic as noted above, but the same is not true of surface knots in $S^{4}$. For instance, the unknotted torus in $S^{4}$ admits two different order 2 periodic symmetries: one in which one of the $S^{1}$ factors of the torus forms a meridian for $\operatorname{Fix}(\tau)$, and another that restricts to the hyperelliptic involution. The former action is free, while the latter fixes four points.
Note that in principle, the intersection of a simply periodic manifold $M$ with the ambient fixed set can be knotted in $M$ (rel. boundary, if applicable). This will be of particular importance to us.
To aid in the exposition of the paper, we make some parametrization conventions. We think of our periodic 2-knot as being contained in the boundary of an upper half-ball $U\subset B^{5}$. We will frequently parametrize $U$ by the coordinates:
$U=\{(x,y,z,w,t)\mid x^{2}+y^{2}+z^{2}+w^{2}+t^{2}\leq C,t\geq 0\}$
for some $C>>1$. When we are doing constructions explicitly in $U\cap S^{4}$, we will refer to coordinates on $U$ simply as $(x,y,z,w)$. We also implicitly choose the standard metric on $U$, scaled appropriately to our choice of coordinates, which we use in our discussion of Morse theory. Submanifolds inside $U$ inherit this metric, and we will frequently differentiate between the gradient of a Morse function on the ambient space and the gradient of its restriction to the submanifold.
Since the action $\tau$ we consider on $B^{5}$ is the linear $\mathbb{Z}/d\mathbb{Z}$ action by rotation about a properly embedded 3-ball with unknotted boundary, we define our parameterization such that in coordinates, $\tau$ fixes the $(0,0,z,w,t)$-hyperplane in $U$.
We will also frequently need to pass between equivariant manifolds and their quotients by $\tau$. Throughout this paper, we denote the quotients of manifolds with overbars; e.g. $Y\subset B^{5}$ becomes $\overline{Y}\subset\overline{B^{5}}$ under $\tau$. (Because $\tau$ is always a linear cyclic action by rotation, the quotient of $B^{5}$ by $\tau$ is always diffeomorphic to $B^{5}$, and the quotient of $S^{4}$ by $\tau$ is always diffeomorphic to $S^{4}$.) To recover the original manifold, we may take the $d$-fold branched cover over the image of the fixed set under the covering map. For this reason, we refer to the image of $\operatorname{Fix}(\tau)$ under the covering map as the branch set and denote it $\operatorname{Branch}(\tau)$.
Finally, there are several points in this paper in which we need to refer to subsets of a manifold with a given Morse function on it. For this we use the following standard convention: if $f$ is a Morse function on a manifold $M$, then:
$M_{c}:=f^{-1}(c)$
for $c\in\mathbb{R}$. We also declare:
$M_{*c}:=f^{-1}(\{a\in\mathbb{R}\mid a*c\})$
where * is an element of ${<,,>,}\{$$\leq$$\geq\}$.
1.2. Equivariant concordance group structure
With these conventions settled, we may formally define the equivariant concordance group. This requires a notion of equivariant connected sum, which in turn requires some extra data. Equivariant connected sum for classical knots was first defined by Sakuma in [Sak86], and we follow the spirit of this construction in our own.
Definition 1.4.
A direction for a $d$-periodic 2-knot $(S,\tau)$ is a choice of orientation on $\operatorname{Fix}(\tau)$ together with a choice of labeling of the two points in $\operatorname{Fix}(\tau)\cap S$ with the labels $p_{+}$ and $p_{-}$.
We call a $d$-periodic 2-knot with a choice of direction a directed $d$-periodic 2-knot.
Definition 1.5.
Let $(S,\tau)$ and $(S^{\prime},\tau^{\prime})$ be two directed $d$-periodic 2-knots. Then we define their equivariant connected sum, denoted $(S\#S^{\prime},\tau\#\tau^{\prime})$ or simply $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$, to be the $d$-periodic 2-knot obtained as follows: First, remove an equivariant neighborhood of $p_{+}$ in $S^{4}$ from the copy of $S^{4}$ containing $S$, and an equivariant neighborhood of $p_{-}^{\prime}$ from the copy of $S^{4}$ containing $S^{\prime}$. Then glue these two copies of $S^{4}$ together along their boundaries such that:
  1. (1)
    the intersections of $\operatorname{Fix}(\tau)$ and $\operatorname{Fix}(\tau^{\prime})$ with the deleted neighborhoods are glued together along their common boundary with opposite orientations (so that the union of $\operatorname{Fix}(\tau)$ with $\operatorname{Fix}(\tau^{\prime})$ has a consistent orientation given by the chosen orientations on each of its components),
  2. (2)
    $\tau$ and $\tau^{\prime}$ agree on the common boundary of the punctured 4-spheres, and
  3. (3)
    the punctured copies of $S$ and $S^{\prime}$ are glued together along their boundaries.
The resulting 2-knot $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$ will then be $d$-periodic under the action given by the actions $\tau$ and $\tau^{\prime}$ on their constituent components.
As in the case with strongly invertible classical knots, equivariant connected sum is not usually commutative; that is, $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$ may not be equivariantly isotopic to $S^{\prime}\#_{\mathbb{Z}/d\mathbb{Z}}S$, where we fix an equivariant map identifying their ambient copies of $S^{4}$ that preserves the orientation on the fixed set. (In the classical dimension, Di Prisa [DP23] has shown that the equivariant concordance group of strongly invertible knots is not itself commutative, and so equivariant connected sum is not even commutative up to concordance. However, we will see that this is not the case for $\mathcal{C}^{2}_{d}$.)
For the following definition, extend the linear order $d$ rotation $\rho$ on $S^{2}$ to $S^{2}\times I$ to $\hat{\rho}$ such that $\hat{\rho}(x,t)=(\rho(x),t)$.
Definition 1.6.
An equivariant concordance of $d$-periodic 2-knots is an embedding $\Psi\colon S^{2}\times I$ into $S^{4}\times I\subset B^{5}$ such that:
$\tau\circ\Psi=\Psi\circ\hat{\rho}$
We will work primarily with $B^{5}$ and think of the concordance taking place inside an embedded copy of $S^{4}\times I$ such that the actions on each $S^{4}$ are the restriction of a global order $d$ rotation on $B^{5}$ fixing a properly embedded unknotted 3-ball. If $(S,\tau)$ and $(S^{\prime},\tau^{\prime})$ are directed, we say that a concordance between them is directed if the orientations on $\operatorname{Fix}(\tau)$ and $\operatorname{Fix}(\tau^{\prime})$ agree. Directed equivariant concordance is then an equivalence relation on the set of directed periodic 2-knots.
Definition 1.7.
Let $C^{2}_{d}$ be the set of directed $d$-periodic 2-knots. The equivariant concordance group $\mathcal{C}^{2}_{d}$ of $d$-periodic 2-knots is the quotient of the set $C^{2}_{d}$ by directed equivariant concordance, with group multiplication given by equivariant connected sum.
It is worth mentioning at this point that the condition of equivariant concordance for 2-knots is well-defined up to equivariant isotopy. For, if $S$ is equivariantly concordant to $S^{\prime}$ and $S^{\prime\prime}$ is equivariantly isotopic to $S$, then we may find an equivariant concordance between $S^{\prime}$ and $S^{\prime\prime}$ by tracing the equivariant isotopy from $S^{\prime\prime}$ to $S$ and then finishing with the concordance from $S$ to $S^{\prime}$.
1.3. Morse theory with vertical boundary
For many constructions in the following sections, we will need a modified version of Morse theory to allow us to work with equivariant manifolds with boundary. The next two subsections are devoted to proving the foundational lemmas we will need for these constructions.
Morse theory with vertical boundary uses Morse functions $f$ on manifolds $M$ with boundary where the associated gradient-like vector field of $f$ is everywhere tangent to $\partial M$. This should not be confused with the notion of a relative Morse function, for which $f$ is constant on connected components of $\partial M$ (e.g. on a cobordism). These two are not mutually exclusive, and one can have a relative Morse function with vertical boundary when $M$ is a manifold with corners. For instance, the height function $f$ on $D^{2}\times I$ given by $f(x,t)=t$ is such a function.
It is important to note that we follow the convention of Kronheimer-Mrowka in [KM07] that the gradient-like vector field is given by $\nabla f$, rather than $\nabla f$. However, we do adapt some of the language from Borodzik-Némethi-Ranicki in [BNR16] to this convention.
Let $\varphi:M\times\mathbb{R}\to M$ be the flow associated to $\nabla f$. Recall that the ascending manifold, denoted $\mathcal{A}(p)$, of a critical point $p$ is the set:
$\mathcal{A}(p):=\{x\in M\mid\lim_{s\to\infty}\varphi(x,s)=p\}$
Similarly, the descending manifold of $p$, denoted $\mathcal{D}(p)$ is the set:
$\mathcal{D}(p):=\{x\in M\mid\lim_{s\to-\infty}\varphi(x,s)=p\}$
We say that an open submanifold $A$ of $M$ is fully vertical with respect to $f$ if for all $x\in A$, $\varphi(x,s)$ is also in $A$ for all $s$ in $\mathbb{R}$. In other words, $A$ is comprised of only trajectories, with no critical points, that are closed under the flow of $\nabla f$. Note that we will use “vertical” as a descriptor for various manifolds throughout the paper; this is not a technical term, and is mostly used in exposition to refer to manifolds that may have critical points, but to which $\xi$ is always tangent.
Definition 1.8.
Let $M$ be a smooth $n$-manifold with boundary, and let $f$ be a Morse function on $M$. Let $p$ be a critical point of $f$ in $\partial M$. Then $p$ is said to be boundary stable if the tangent space $T_{p}\mathcal{A}(p)$ is contained entirely within $T_{p}\partial M$. Else, $p$ is said to be boundary unstable.
We say that the index $\text{ind}(p)$ of a boundary critical point is $\dim\mathcal{D}(p)$. This definition agrees with the usual index of an interior critical point. However, note that with this definition, there can be no boundary stable critical points of index 0, nor boundary unstable critical points of index $n$.
[…]
Figure 1. Left: a boundary stable critical point. Right: a boundary unstable critical point. Ascending and descending manifolds are shaded or shown as arrows according to their dimension.
Definition 1.9 ([BNR16]).
An $n$-dimensional index $k$ right half-handle $R^{k}$ is given by:
$R^{k}:=D^{k}\times D^{n-k}$
When the index is clear, we simply write $R$. The boundary $\partial R$ is subdivided into three pieces. One piece is the attaching region $A^{k}$, given by:
$A^{k}\colon=\partial D^{k}\times D^{n-k}$
The remainder of the boundary is topologically a copy of $D^{k}\times S^{n-k-1}$. Decompose $S^{n-k-1}$ into two copies $D_{h}^{n-k-1}$ and $D_{v}^{n-k-1}$ of $D^{n-k-1}$ glued along their boundaries. Then the other two pieces of $\partial R^{k}$ we call the horizontal and vertical portions of the boundary, denoted $(\partial R)_{h}$ and $(\partial R)_{v}$, respectively, given by:
$(\partial R)_{h}=D^{k}\times D_{h}^{n-k-1}$
$(\partial R)_{v}=D^{k}\times D_{v}^{n-k-1}$
This definition should be interpreted as follows. Suppose $M$ is an $n$-manifold with boundary and fix a relative Morse function $f$ on $M$ such that $f$ is constant on some connected component of $\partial M$. Then an index $k$ right half-handle $R^{k}$ can be attached to that portion of $\partial M$ along $A^{k}$. However, we do this attachment so that $(\partial R)_{h}$ is horizontal with respect to $f$, and $(\partial R)_{v}$ is vertical with respect to $f$. Thus a right half-handle attachment changes both the topology of $\partial M$ and that of $M$ itself by attaching an index $k$ handle to each.
There is also a notion of left half-handles, but we will not use them in this paper. See [BNR16] for more details.
Figure 2. A 3-dimensional index 1 right half-handle attachment. Note the shading to denote the filling of the interior manifold after the boundary handle attachment.
Borodzik-Némethi-Ranicki showed, in analogy with classical Morse theory, that passing a boundary unstable critical point induces a right half-handle attachment. We will use this notion to build Seifert manifolds by hand by declaring that whenever we pass a critical point of the boundary, we attach a right half-handle to the Seifert manifold.
1.4. Equivariant Morse theory for finite group actions
We begin with a few remarks about equivariant Morse theory in general. For our purposes, and in much of the literature of equivariant Morse theory, the action of $\mathbb{Z}/d\mathbb{Z}$ on the codomain $\mathbb{R}$ is assumed to be trivial. Thus when we say that a Morse function $f$ is equivariant, we mean that $f\circ\tau=f$ where $\tau$ is the action on the domain.
Equivariant Morse theory for compact manifolds was studied by Wasserman in [Was69], building on work of Bott [Bot54] and Palais [Pal63]. Bao and Lawson [BL25] recently gave an extension of this result that shows genericity of Morse-Smale functions among pairs of equivariant functions and equivariant metrics. We will use this result implicitly throughout this subsection. Wasserman also showed, for compact Lie groups $G$, that the correspondence between critical points of a Morse function and handle attachment carries over to the $G$-equivariant setting by defining a critical $G$-orbit. In the case of our finite order rotation actions, equivariant handles always appear as either a single handle that is fixed setwise by $\tau$, or as a $d$-tuple of handles that are permuted by $\tau$.
It should be noted that in general, equivariant differential topology is much more restrictive than the non-equivariant setting. Transversality results often fail equivariantly. However, with sufficient conditions on the group action, the fixed set, and the manifold acted on, many standard results in Morse theory can be recovered in the equivariant setting. For instance, in the context of equivariant 4-manifold topology, Meier and Scott were able to prove an equivariant version of the Laudenbach-Poénaru theorem [MS25a], which enabled them to study equivariant trisections and bridge trisected surfaces [MS25b]. We will not use these results explicitly, but we mention them due to their close relationship to our study of equivariant 2-knots.
Definition 1.10.
Let $M^{k}$ be smoothly embedded in $B^{m}$ for $m-k\geq 2$ and equivariant with respect to the standard $\mathbb{Z}/d\mathbb{Z}$ action $\tau$ by rotation fixing a properly embedded copy of $B^{m-1}$. Let $f$ be an equivariant Morse function on $B^{m}$. For each critical point $p$ of $f$ restricted to $M$, let $\text{ind}(p)$ denote its index. We say that $M$ is in ordered equivariant Morse position with respect to $f$ if:
  1. (1)
    the critical points of $f|_{M}$ are isolated,
  2. (2)
    the Hessian of $f|_{M}$ at each critical point is nondegenerate,
  3. (3)
    every critical point of $f|_{M}$ is either fixed by $\tau$ or sent to another critical point of the same index, and
  4. (4)
    for all critical points $p_{i}$ and $p_{j}$, $f(p_{i})<f(p_{j})$ implies that $\text{ind}(p_{i})\leq\text{ind}(p_{j})$.
Recall that the purpose of our study of equivariant Morse theory is to build equivariant Seifert manifolds. In order to do this, we will need the boundary manifolds to be in ordered equivariant Morse position. For completeness, we prove a stronger statement that applies generally to embedded manifolds with finite order actions fixing a codimension 2 submanifold. This requires several steps, which are outlined in the following lemmas. The formal statement we wish to prove is:
Proposition 1.11.
Let $M$ be an $n$-manifold, possibly with boundary, embedded in $B^{m}$ where $m-n\geq 2$. Fix an equivariant Morse function $f$ on $B^{m}$ with two boundary critical points (one of index 0 and the other of index $m$), and let $\psi\colon M\to B^{m}$ be a $\tau$-equivariant embedding such that $A:=\operatorname{Fix}(\tau)\cap\psi(M)$ has at most one maximum and one minimum, with no other critical points with respect to $f$. Then $\psi$ is equivariantly isotopic to an embedding $\psi^{\prime}$ such that $\psi^{\prime}(M)$ is in ordered equivariant Morse position with respect to $f$, and $A^{\prime}:=\operatorname{Fix}(\tau)\cap\psi^{\prime}(M)$ has the same number and type of critical points as $A$.
The first step towards proving this is to describe when it is possible to perturb a Morse function on a manifold embedded in $\mathbb{R}^{n}$ to rearrange critical points in the complement of a fully vertical submanifold. This is essentially an adaptation of Theorem 4.1 of [BP16], and our proof follows theirs closely, but we here we are careful to keep the perturbation of the Morse function away from the fully vertical submanifold. We also only move one critical point of the Morse function restricted to the submanifold, deviating from [BP16] and other sources (e.g. [Mil65]) that contain this lemma that simultaneously rearrange a pair.
Lemma 1.12.
Let $M$ be a compact $n$-manifold, possibly with boundary or corners, embedded in $W\cong B^{m-1}\times[0,1]$. Let $A$ be a properly embedded codimension 2 submanifold of $M$. Let $f\colon W\to[0,1]$ be a relative Morse function such that:
  1. (1)
    $f$ has no critical points on $W$ ($f$ is a submersion),
  2. (2)
    $f|_{M}$ has exactly two (possibly boundary) critical points, denoted $p_{k}$ and $p_{l}$ of indices $k$ and $l$ respectively, and
  3. (3)
    AA is fully vertical with respect to $f$.
Suppose that $k<l$ but $f(p_{k})>f(p_{l})$. Then there exists a Morse function $g$ on $W$ that also has no critical points on $W$, is homotopic to $f$ through submersions of $W$, and with $g(p_{k})<g(p_{l})$. Furthermore, $g$ can be chosen such that $f$ and $g$ agree on a neighborhood of $A$.
Proof.
For $p\in\{p_{k},p_{l}\}$, let $\mathcal{M}(p)=\mathcal{A}(p)\cup\{p\}\cup\mathcal{D}(p)$. Then since $k<l$, a codimension counting argument shows that $\mathcal{M}(p_{k})\cap\mathcal{M}(p_{l})$ is generically empty. Furthermore, since $A$ was assumed to be fully vertical, both $\mathcal{M}(p_{k})$ and $\mathcal{M}(p_{l})$ are disjoint from $A$.
Choose a neighborhood $N_{0}$ of $\mathcal{A}(p_{k})\cap(B^{m-1}\times\{0\})$ that is disjoint from a neighborhood of $\mathcal{A}(p_{l})\cap(B^{m-1}\times\{0\})$ and from a neighborhood of $A\cap(B^{m-1}\times\{0\})$. Let $\mu\colon B^{m-1}\times\{0\}\to\mathbb{R}$ be a smooth bump function supported on $N_{0}$ such that $\mu|_{\mathcal{A}(p_{k})\cap(B^{m-1}\times\{0\})}\equiv 1$. Then $\mu$ can be extended to a map $\hat{\mu}\colon W\to\mathbb{R}$ by extending along the flow of a gradient-like vector field of $f$. That is, if $\varphi\colon W\times\mathbb{R}\to W$ is the flow generated by $\nabla f$, then:
$\hat{\mu}(x)=\mu\left(\lim_{s\to-\infty}\varphi(x,s)\right)$
In particular, $\hat{\mu}$ is constant along trajectories of $f$. Let $N$ be the extension of $N_{0}$ to $W$ under this flow, so that $N$ is a neighborhood of $\mathcal{A}(p_{k})$ that is disjoint from neighborhoods of $\mathcal{A}(p_{l})$ and $A$.
Choose $a\in[0,1]$. This will be our target value for our new Morse function at $p_{k}$. We construct the new Morse function by rearranging the heights of points in $W$ based on their $\hat{\mu}-$value. This is done by choosing a function $\Upsilon\colon[0,1]\times[0,1]\to[0,1]$ satisfying:
  1. (1)
    $\frac{\partial\Upsilon}{\partial r}(r,u)>0$ for all $r$ and $u$,
  2. (2)
    $\Upsilon(f(p_{k}),1)=a$,
  3. (3)
    $\Upsilon(r,u)=r$ for $r$ in small neighborhoods of 0 and 1 and all $u$, and also for $u$ in a small neighborhood of 0
  4. (4)
    $\frac{\partial\Upsilon}{\partial r}(r,1)=1$ for $r$ in a neighborhood of $f(p_{k})$, and $\frac{\partial\Upsilon}{\partial r}(r,0)=1$ for $r$ in a neighborhood of $f(p_{l})$.
Given such a function $\Upsilon$, we may define our new Morse function by:
$g(w)=\Upsilon(f(w),\hat{\mu}(w))$
Condition (1) ensures that $g$ has no critical points on $W$. Condition (2) is simply the statement that $g(p_{k})=a$. Condition (3) requires that $f$ and $g$ agree on the horizontal boundary of $\partial W$ and away from $N$. Finally, condition (4) ensures that $g$ is essentially a linear shift of $f$ near the critical points $p_{k}$ and $p_{l}$, which preserves their type (interior vs. boundary) and index.
We see that since $g$ and $f$ agree away from $N$, they in particular agree on a neighborhood of $A$. Through a continuous family of choices for values of $a$ terminating with $a<f(p_{l})$, we find a family of Morse functions with no critical points on $W$ interpolating between $f$ and a Morse function $g$ with $g(p_{k})<g(p_{l})$, and with $A$ remaining fully vertical throughout. ∎
It is not enough to find new ambient Morse functions that rearrange critical points, however. For Proposition 1.11, we need these perturbations to be realized by an isotopy of the embedded manifold. We thank Mark Powell for his help with the following lemma.
Lemma 1.13.
Let $M$ be a compact $n$-manifold, $W$ be a compact $m$-dimensional cobordism, and let $\psi$ be a smooth embedding of $M$ in $W$. Suppose that $f$ and $g$ are two relative Morse functions on $W$ that each have no critical points on $W$ and are homotopic through submersions of $W$. Then there exists an isotopy $\Psi\colon M\times[0,1]\to W$ and an orientation-preserving diffeomorphism $h\colon\mathbb{R}\to\mathbb{R}$ such that $\Psi(x,0)=\psi(x)$ and $\Psi(x,1)$ satisfies:
$f\circ\Psi(x,1)=h\circ g\circ\psi(x)$
Note that, as in Lemma 1.12, $f$ and $g$ may have critical points when restricted to $\psi(M)$, but cannot have any on $W$ itself. We remind the reader of the definition of a stable map to aid in the proof. For more on stable maps, the author recommends [GG73].
Definition 1.14.
Let $f\in C^{\infty}(X,Y)$ for $X$ and $Y$ smooth manifolds. Then $f^{\prime}\in C^{\infty}(X,Y)$ is said to be equivalent (in the sense of stable maps) to $f$ if there exist orientation-preserving diffeomorphisms $\phi\colon X\to X$ and $\eta\colon Y\to Y$ such that the following diagram commutes:
[…]
A map $f$ is said to be stable if there exists a neighborhood $N(f)\subset C^{\infty}(X,Y)$ such that all $f^{\prime}\in N(f)$ are equivalent to $f$.
Proof of Lemma 1.13.
Let $F\colon W\times[0,1]\to\mathbb{R}$ be a homotopy through relative Morse functions with no critical points such that $F(w,0)=f(w)$ and $F(w,1)=g(w)$. Then $F(w,t)$ is Morse for all $t$. Morse functions are stable so long as they do not take the same value on any two critical points [GG73, Ch. III Prop. 2.2], so for each $t$, there exists a neighborhood N(F(,t))N(F(-,t)) in which all functions are equivalent to F(,t)F(-,t). We therefore have that all the Morse functions F(,t)F(-,t) are equivalent to each other, and so for each $t$ we have maps $\phi_{t}$ and $\eta_{t}$ such that the following diagram commutes:
[…]
So if we let $\Psi(x,t)=\phi_{t}^{-1}(\psi(x))$, then our diagram becomes:
[…]
And so we see that:
$f\circ\Psi(x,1)=\eta_{1}^{-1}\circ F(\psi(x),1)=\eta_{1}^{-1}\circ g\circ\psi(x)$
This completes the proof. ∎
For our purposes, we need not worry about the function $\eta_{1}^{-1}$ when finding an isotopy interpolating between $f$ and $g$, as orientation preserving diffeomorphisms of $\mathbb{R}$ are all isotopic to the identity and merely contribute a rescaling of our Morse function that we can undo later.
Remark 1.15.
From the proof of stability for Morse functions with distinct critical values (see [GG73, Ch. III Prop. 2.2]), one sees that the functions $\phi_{t}$ are constructed so that they are supported on neighborhoods of the subsets of $W$ on which F(,t)F(-,t) differs from $f$. Thus the image of the support of the isotopy $\Psi$ can be taken to be contained in a neighborhood of the support of fgf-g.
We can now prove the existence of equivariant isotopies to ordered equivariant Morse position, given a fixed set with only two critical points.
Proof of Proposition 1.11.
Recall that we begin with an equivariant Morse function $f$ on $B^{m}$, and that $A=\psi(M)\cap\operatorname{Fix}(\tau)$. By assumption, $f|_{A}$ has at most two critical points; denote these $p_{+}$ and $p_{-}$ according to whether they are maxima or minima. If $p_{+}$ exists, choose an equivariant neighborhood of $p_{+}$ and perturb $\psi$ to isotope it “up” along a path in $\operatorname{Fix}(\tau)$ such that $f(p_{+})>f(x)$ for all $x\in M$, using a bump function to smooth the corners. Such a path exists because the fixed set of the standard rotation action on $B^{m}$ is a copy of $B^{m-2}$, and thus is connected. Similarly, if $p_{-}$ exists, perturb $\psi$ to isotope an equivariant neighborhood of $p_{-}$ “down” so that $f(p_{-})<f(x)$ for all $x$ in $M$ and smooth the corners with a bump function. This procedure guarantees that the lowermost index 0 critical point and the uppermost index $n$ critical point of $M$ are both in $A$.
Let $\overline{M}$ be the quotient of $M$ by the pullback of $\tau$, and consider the quotient embedding $\overline{\psi}\colon\overline{M}\to\overline{B^{m}}$. (Note that $\overline{B^{m}}$ is still diffeomorphic to $B^{m}$; we use the symbol $\overline{B^{m}}$ to emphasize that we are working in the quotient manifold.) Since $f$ is equivariant, the image of $p_{+}$ and $p_{-}$ under the quotient map retain their maximal and minimal positions with respect to the induced Morse function $\overline{f}$ on $\overline{B^{m}}$.
Let $a_{\pm}=\overline{f}(\overline{p}_{\pm})$, and let $\overline{W}=\overline{f}^{-1}([a_{-}+\epsilon,a_{+}-\epsilon])$ for $\epsilon$ small enough that the only two critical points of $\overline{f}|_{\overline{\psi}(\overline{M})}$ outside of $\overline{W}$ are $a_{+}$ and $a_{-}$. Then $\overline{M}\cap\overline{W}$ is a cobordism between two unknotted spheres (or disks, if $\partial M\neq\emptyset$) embedded in $B^{m-1}\times[0,1]$. The intersection of $\overline{\psi}(\overline{M})\cap W$ with $\operatorname{Branch}(\tau)$ is a copy of $S^{n-3}\times I$ embedded vertically with respect to $\overline{f}$. Denote this intersection $\overline{A}$, and note that because $\overline{f}$ is equivariant, $\overline{A}$ is fully vertical with respect to $\overline{f}$.
Since $f$ has no critical points on $W$, we may use Lemma 1.12 to successively perturb $f$ to obtain Morse functions $f_{i}$ that rearrange pairs of critical points in the complement of $\operatorname{Branch}(\tau)$. The perturbations are supported on neighborhoods of descending manifolds of lower index critical points, and each $f_{i}$ is related to $f_{i+1}$ by a homotopy through submersions. Hence by Lemma 1.13, we may find an isotopy $\overline{\Psi}\colon\overline{M}\times[0,1]\to\overline{W}$ realizing this rearrangement, with $\overline{\Psi}(-,0)=\psi$. This can then be extended to an ambient isotopy (which, by abuse of notation, we also call $\overline{\Psi}$) by the Isotopy Extension Theorem. As we noted in Remark 1.15, since each perturbation of the Morse function was disjoint from $\operatorname{Branch}(\tau)$, $\overline{\Psi}$ can be taken to be supported in the complement of a neighborhood of $\operatorname{Branch}(\tau)$.
It remains to lift these rearrangements back to the cover. For this, we use Schwarz’s resolution of the Isotopy Lifting Conjecture [Sch80], which implies that isotopies in the quotient of a manifold by a compact Lie group can be lifted to isotopies of the manifold itself if they are supported away from the image of the fixed set. Thus we may lift $\overline{\Psi}$ to an isotopy $\Psi\colon M\times[0,1]\to B^{m}$ such that $\Psi(-,0)=\psi$ and the image $\Psi(M,1)$ is in ordered equivariant Morse position, again extended to an ambient isotopy. ∎
2. Equivariant Seifert solids
We now proceed to our construction of equivariant Seifert 3-manifolds. Our proof adapts a technique of Yanagawa [Yan69].
Proposition 2.1.
Let $S$ be a periodic 2-knot in $S^{4}$. Then $S$ bounds a $\tau$-equivariant Seifert 3-manifold $Y\subset S^{4}$ such that $\operatorname{Fix}(\tau)\cap Y$ consists of an arc $A$ that is null-homologous in $Y$ rel. boundary.
The remainder of this section is devoted to the construction of $Y$.
Proof.
Fix an equivariant Morse function $f$ on $S^{4}$ such that its restriction to a ball around $S$ has exactly two critical points. By Proposition 1.11, we may assume that $S$ is in ordered equivariant Morse position with respect to $f$. Note that without this proposition, we would be unable to assume that the ambient Morse function induces a self-indexing Morse function on $S$, which is crucial to our construction.
It will be convenient to make use of our established conventions for coordinates for this neighborhood of $S$, as described in the introduction. In particular, we use coordinates $(x,y,z,w)$ such that $f$ is projection onto the $w$-axis and $\tau$ fixes the $(z,w)$-subspace.
First, recall that in the proof of Proposition 1.11, the critical points of the fixed set of the embedded manifold $M$ were first isotoped to be the highest maximum and lowest minimum with respect to the ambient Morse function. In the case of a 2-sphere, these are actually the only two fixed points, and so by assuming that $S$ is in ordered equivariant Morse position, we assume that the lower fixed point $p_{-}$ is the global minimum of $f|_{S}$ and that the higher, $p_{+}$, is its global maximum. In particular, the descending manifolds of each of the index 1 critical points do not contain $p_{+}$ or $p_{-}$, and so do not intersect $\operatorname{Fix}(\tau)$. Note further that apart from $p_{+}$ and $p_{-}$, all the other critical points of $K$ must come in $d$-tuples that are permuted by $\tau$. Up to translation and stretching of our coordinates, we may now assume that all the index $k$ critical values of $f|_{S}$ occur within $(k,k+\epsilon)$.
We remind the reader of our notation. By $S_{w}$, we mean the preimage of $w$ under $f$ intersected with $S$, thought of as living in the whole 3-dimensional preimage of $w$. We call this preimage a $w$-slice. In this perspective, $\tau$ appears in each $w$-slice as rotation about the $z$-axis. We may construct $Y$ by specifying $Y_{w}$, defined analogously, for each $S_{w}$. In particular, for regular values $w$ of $f$, $S_{w}$ is a periodic link, and $Y_{w}$ is a Seifert surface for that link intersecting the $z$-axis in a single point. A smoothly varying family of such surfaces then stack together to form $Y$ with the desired fixed set. We will regularly switch between this perspective and explicit handle attachments to $Y$ throughout the construction. A schematic of the various handle attachments in the construction of $Y$ is shown in Figure 3 for the reader to follow along as we proceed through the construction.
[…]
Figure 3. A schematic of $Y$ showing handle attachments.
$\boldsymbol{f=0.}$ Since $S$ is in ordered equivariant Morse position, we may consider all of its 0-handles first. These will be a collection of unlinked, unknotted disks. For each of these disks, we begin our construction of $Y$ by attaching a 3-dimensional index 0 right half-handle. Since all these attachments occurred in $f^{-1}(0,\epsilon)$, we see that $S_{\nicefrac{{1}}{{2}}}$ is a collection of unlinked, unknotted circles such that one of them links the $z$-axis and the rest are permuted by $\tau$, and $Y_{\nicefrac{{1}}{{2}}}$ appears as a collection of standardly embedded disks bounded by each component of $S_{\nicefrac{{1}}{{2}}}$. This is illustrated on the left of Figure 4.
$\boldsymbol{f=\frac{1}{2}.}$ Now for each symmetric $d$-tuple of index 1 critical points, we consider the corresponding descending manifolds. These consist of collections of arcs passing through the index 1 critical points, possibly linking components of $S_{w}$ for $w$ below the critical value, and eventually intersecting those unknotted components. Up to vertical isotopy, these intersections can be chosen to occur in $f^{-1}(\frac{1}{2},\frac{1}{2}+\epsilon)$. A schematic of this is shown on the right in Figure 4.
Figure 4. Left: the central 0-handle of $Y$, shown in multiple $w$-slices. Right: index 1 trajectories of $S$ linking slices of $S$, which would pierce the disks on the left.
Because the descending manifolds may link the components of $S_{w}$ for $\frac{1}{2}+\epsilon<w<1$, we cannot simply continue the existing 0-handles of $Y$ upwards, as they would then intersect the descending manifolds. To avoid this, we first equivariantly perturb $S$ to introduce new intersections of the descending manifolds with disks bounded by $S_{\nicefrac{{1}}{{2}}}$ as necessary so that each descending arc intersects the disks in pairs of oppositely signed points. We then attach equivariant 3-dimensional interior 1-handles to $Y$ whose corresponding interior critical points have critical values in $(\frac{1}{2}-\epsilon,\frac{1}{2})$ and whose attaching spheres are vertical translates of the intersections of the index 1 boundary trajectories with the disks bounded by $S_{\nicefrac{{1}}{{2}}}$. In each $w$-slice for $w$ above $\frac{1}{2}$, we therefore see equivariant tubes attached to the disks bounded by $S_{w}$ so that the descending manifolds of the index 1 boundary critical points are engulfed by these tubes and thus do not intersect our Seifert surface in those cross-sections. This is illustrated in Figure 5.
Figure 5. A diagram of $Y_{\nicefrac{{3}}{{2}}}$, showing $S_{\nicefrac{{3}}{{2}}}$ as a periodic unlink bounding a symmetric surface. Note that the descending manifolds of the index 1 boundary critical points, depicted as bands, are prevented from intersecting the obvious disks bounded by each component by introducing tubes for them to pass through.
$\boldsymbol{f=1.}$ We may now consider each index 1 critical point of $S$ as a boundary unstable index critical 1 point of $Y$, and thus for each 1-handle of $S$, we get one right half-handle attachment for $Y$. After performing the above procedure for all $d$-tuples of 1-handles in $S$, the cross-section $S_{\nicefrac{{3}}{{2}}}$ will consist again of a collection of unknotted components, one of which is a meridian for the $z$-axis. The corresponding surface $Y_{\nicefrac{{3}}{{2}}}$ will be a possibly disconnected equivariant surface of unknown genus with boundary $S_{\nicefrac{{3}}{{2}}}$.
$\boldsymbol{f=\frac{3}{2}.}$ At this point, $S_{>\nicefrac{{3}}{{2}}}$ consists only of disks, but $\partial(Y_{\leq\nicefrac{{3}}{{2}}})$ is a potentially complicated equivariant surface. We would like to cap it off with equivariant 3-handles, but to do this, we must first surger it to a collection of disks. Therefore we now proceed to find a collection of equivariant compressing curves on $Y_{\nicefrac{{3}}{{2}}}$ such that the compressions do not introduce any new intersections with $\operatorname{Fix}(\tau)$.
Let $\{J_{i}\}_{i\in I}$ be the set of connected components of $S_{\nicefrac{{3}}{{2}}}$, with $J_{0}$ the component that links the $z$-axis. Then the components $\{J_{i}\}_{i\in I}$ bound a pairwise disjoint collection of maximum disks $\{D_{i}\}_{i\in I}$ of $S$ such that only $D_{0}$ intersects $\operatorname{Fix}(\tau)$, and it does so transversely and exactly once. These $D_{i}$ are themselves equivariant. We may therefore equivariantly isotope copies of these disks into $f^{-1}(\frac{3}{2})$ by flowing along $\nabla f$. Call these isotoped disks $D_{i}^{\prime}$. Then equivariant $\epsilon$-neighborhoods of the disks $D_{i}^{\prime}$ for $\epsilon$ sufficiently small have boundaries that form splitting spheres $V_{i}$ separating each component of $S_{\nicefrac{{3}}{{2}}}$ from each other, and separating all $J_{i}$ except $J_{0}$ from the $z$-axis in the $\frac{3}{2}$-slice.
We can now make a standard “innermost circles” argument to find equivariant compressing disks for $Y_{\nicefrac{{3}}{{2}}}$. Consider the intersections of each splitting sphere $V_{i}$ with $Y_{\nicefrac{{3}}{{2}}}$. These intersections form nested concentric circles on each sphere $V_{i}$. By compressing along these concentric disks starting with the innermost circle of intersection and working outward, we can systematically compress $Y_{\nicefrac{{3}}{{2}}}$ without introducing new intersections with the splitting spheres, as shown in Figure 6. Since these spheres are disjoint from the $z$-axis, and lie entirely in a single $w$-slice, the compressing disks are also disjoint from $\operatorname{Fix}(\tau)$, and so compressing along them does not introduce any new intersections with $\operatorname{Fix}(\tau)$. Actually performing the compressions is done by the equivariant attachment of interior 2-handles to $Y_{\leq\nicefrac{{3}}{{2}}}$, where the core disks of the handles are taken to be the compressing disks themselves. We declare that these handle attachments occur within $f^{-1}(\frac{3}{2},\frac{3}{2}+\epsilon)$.
[…]
Figure 6. “Innermost circles” compression of $Y_{\nicefrac{{3}}{{2}}}$ using compressing disks obtained via the boundary sphere $V_{i}$ of a tubular neighborhood of one of the disks $D^{\prime}_{i}$.
$\boldsymbol{f=2.}$ After all such compressions are performed, $S_{\nicefrac{{3}}{{2}}+\epsilon}$ is unchanged, but $Y_{\nicefrac{{3}}{{2}}+\epsilon}$ now consists of an equivariant collection of disjoint disks and closed components, still intersecting the $z$-axis in a single point contained inside the central disk. We further compress the closed components until they are spheres, and then fill them in with interior 3-handles. This also does not change $S_{\nicefrac{{3}}{{2}}+\epsilon}$, but surgers $Y_{\nicefrac{{3}}{{2}}+\epsilon}$ to a collection of disks that are equivariantly isotopic to the maximum disks of $S$. Thus we may fill in the remainder of $S$ by declaring $Y_{\geq\nicefrac{{3}}{{2}}+\epsilon}$ to be trace of this isotopy. This gives $Y$, and its intersection $A$ with $\operatorname{Fix}(\tau)$ is precisely its intersection with the $w$-axis, which is a single arc as desired.
It remains to show that the intersection $A$ of $Y$ with $\operatorname{Fix}(\tau)$ is null-homologous in $Y$ rel. boundary. Indeed, $A$ does not intersect the belt spheres of any 1-handles of $Y$, as all of the 1-handles were chosen to miss $\operatorname{Fix}(\tau)$. Thus $A$ is 0 in the cellular homology rel. boundary induced by the handle decomposition. ∎
We will return to our computation of the concordance group of equivariant 2-knots after a brief detour to define an invariant that will become our obstruction to equivariant sliceness.
3. An extension of the Arf invariant
3.1. Definition and related notions
The Arf invariant of a quadratic form was introduced by Arf in [Arf41] as a purely algebraic tool. Robertello gave the first definition of the Arf invariant of a knot in $S^{3}$ in [Rob65]. His paper describes three invariants of knots in $S^{3}$, then proves they are all equivalent. For our purposes, it will be helpful to recall all three. Recall that a surface $F$, possibly with boundary, properly embedded in a 4-manifold $X$ is called characteristic if $F]\in H_{2}(X,\partial X;\mathbb{Z}/2\mathbb{Z})$ is dual to the second Stiefel-Whitney class of $X$. Equivalently, for all $x\in H_{2}(X,\partial X;\mathbb{Z}/2\mathbb{Z})$, we have that $F]\cdot x\equiv x\cdot x$ mod 2, where $\cdot$ denotes intersection number rel. boundary.
Definitions 3.1 (Robertello).
Let $K$ be a knot in $S^{3}$.
  1. (1)
    Let $F$ be a Seifert surface for $K$ of genus $g$. Choose a symplectic basis for $H_{1}(F;\mathbb{Z}/2\mathbb{Z})$ and let $I_{F}\colon H_{1}(F;\mathbb{Z}/2\mathbb{Z})\times H_{1}(F;\mathbb{Z}/2\mathbb{Z})\to\mathbb{Z}/2\mathbb{Z}$ be the associated Seifert form, considered as a $2g\times 2g$ matrix over $\mathbb{Z}/2\mathbb{Z}$ in the chosen basis. Let $q_{F}\colon H_{1}(F;\mathbb{Z}/2\mathbb{Z})\to\mathbb{Z}/2\mathbb{Z}$ be given by:
    $q_{F}(v)=I_{F}(v,v)$
    Then $q_{F}$ is a quadratic form. Arf invariant of $K$ is the Arf invariant of $q_{F}$.
  2. (2)
    Let $X$ be a smooth 4-manifold with boundary $S^{3}$ such that $K$ bounds a smooth, properly embedded characteristic disk $D$ into $X$. Define the quantity $\varphi(K)\in\mathbb{Z}/2\mathbb{Z}$ by:
    $\varphi(K):=\frac{\sigma(X)-D\cdot D}{8}\mod 2$
    where $D\cdot D$ here represents the normal Euler number of $D$ relative to the Seifert framing of $K$.
  3. (3)
    We say $K$ and $J$ are band pass equivalent if there exists a sequence of planar isotopies, Reidemeister moves, and band passes taking a diagram of $K$ to a diagram of $J$. Band pass equivalence is an equivalence relation on the set of knots in $S^{3}$, and divides knots into two classes: those band pass equivalent to the unknot, and those band pass equivalent to the trefoil. We see therefore that the band pass equivalence class is an invariant of $K$.
Theorem 3.2 (Robertello).
The three definitions given above are all equivalent to each other. Specifically, $K$ has Arf invariant 0 if and only if $\varphi(K)$ vanishes, and equivalently, if and only if it is band pass equivalent to the unknot.
Most recent treatments of the Arf invariant of knots use the definition (1); for instance, this is the definition given in [Sav12], Chapter 9. However, there are also several extensions and related invariants in 3- and 4-manifold topology that build on these equivalent definitions. For instance, the Arf invariant was successfully extended to knots in integer homology 3-spheres, where it interacts with the Rokhlin invariant of the ambient 3-manifold, see [Gor75]. There is also a notion of the Arf invariant of a properly embedded characteristic surface in a 4-manifold. A modern discussion of this invariant can be found in [Sav12]. However, most comprehensive account of all of these invariants and their relation to each other is given by Klug [Klu21], where he proves the following for a characteristic surface $F$ properly embedded in a smooth 4-manifold $X$ with $\partial F=K$ and $\partial X$ an integer homology 3-sphere:
$\operatorname{Arf}(F)+\operatorname{Arf}(K)\equiv\frac{\sigma(X)-F\cdot F}{8}+\mu(\partial X)\mod 2.$
Here $\mu$ denotes the Rokhlin invariant.
Below we give a generalization of the Arf invariant to null-homologous knots in arbitrary oriented 3-manifolds based primarily on definitions (2) and (3). Throughout, all knots are null-homologous in their ambient 3-manifolds and therefore have well-defined Seifert longitudes. For any embedded surface $F$ in a 4-manifold bounded by such a knot, we denote by $F\cdot F$ the normal Euler number of $F$ relative to the Seifert framing on the boundary.
Definition 3.3.
Let $K$ be a closed, null-homologous knot embedded in a 3-manifold $Y$. Let $X$ be a spin cobordism between $Y$ and $S^{3}$. Let $C$ be a properly embedded characteristic annulus in $X$ with normal Euler number $C\cdot C\equiv 0$ mod 16 such that $\partial C=K\sqcup J$, where $J$ is a knot in $S^{3}$. We define the Arf invariant of $K$ in $Y$, denoted $\operatorname{Arf}(K,Y)$, to be:
$\operatorname{Arf}(K,Y):=\operatorname{Arf}(J)\in\mathbb{Z}/2\mathbb{Z}$
To aid in our discussion of this invariant, we will call an annulus satisfying the conditions in Definition 3.3 an Arf annulus.
3.2. Well-definedness
To show that $\operatorname{Arf}(K,Y)$ is well-defined, it is helpful to establish the existence of spin fillings with Arf annuli for all null-homologous knots in 3-manifolds.
Proposition 3.4.
For all null-homologous $K\subset Y$, there exists a spin cobordism $X$ between $Y$ and $S^{3}$ and an Arf annulus $C$ bounded on one end by $K$.
Proof.
By a result of Kaplan [Kap79], for any spin structure on $Y$ there always exists a spin cobordism between $Y$ and $S^{3}$ consisting of only 2-handles, and so we may interpret this cobordism as the result of spin surgery on a collection of framed curves in $Y$. Since $K$ is null-homologous, we may assume that $K$ is also null-homologous in the complement of the surgery curves, and in particular that $K$ bounds a surface $F$ in $Y$ disjoint from all the surgery curves. (To ensure this, we can start with any surface bounded by $F$ and introduce more genus to remove intersections with the surgery curves.) Thus we can obtain $C$ by tracing $K$ through successive cross-sections of the surgery: if $X$ is given by $Y\times I\cup\{2\text{-handles}\}$, then $C$ is given by $K\times I$. Furthermore, since none of the surgery curves intersect $F$, the product $F\times I$ embeds properly in $X$, and so the Seifert framings on $K$ and $J$ extend over $C$. We conclude that $C$ is 0-framed rel. boundary. ∎
[…]
Figure 7. Schematic of an Arf annulus.
We also state a corollary of Theorem 3.2 that applies only to spin manifolds.
Corollary 3.5.
Let $K$ be a knot in $S^{3}$. Let $X$ be a spin filling of $S^{3}$ and let $D$ be a characteristic disk in $X$ bounded by $K$. Then:
$\operatorname{Arf}(K)\equiv\frac{D\cdot D}{8}\mod 2.$
Proof.
Rokhlin’s theorem [Roh52] states that spin 4-manifolds have vanishing signature mod 16. Thus by the equivalence of definitions (2) and (3) in Definitions 3.1, we obtain the result. ∎
With this established, well-definedness follows almost immediately. In the following proposition, for a manifold or knot $M$, M-M indicates $M$ with the opposite orientation.
Proposition 3.6.
Let $C_{1},C_{2}$ be two Arf annuli in different spin cobordisms $X_{1},X_{2}$ from $Y$ to $S^{3}$ as in Proposition 3.4, and let $J_{1}$, $J_{2}$ be their other boundary components, respectively. Then $\operatorname{Arf}(J_{1})=\operatorname{Arf}(J_{2})$.
Proof.
Without loss of generality, assume for the purposes of contradiction that $\operatorname{Arf}(J_{1})=1$ and $\operatorname{Arf}(J_{2})=0$. Let $X$ be obtained by gluing $X_{1}$ and $X_{2}$ along $Y$, and let $C$ be the annulus obtained by gluing $C_{1}$ and $C_{2}$ along $K$. Then $X$ is a spin cobordism from $S^{3}$ to $S^{3}$, and $C$ is an Arf annulus from $J_{1}$ to $J_{2}$. Since $\operatorname{Arf}(J_{2})=0$, we have $\operatorname{Arf}(-J_{2})$ = 0 as well. Thus by Proposition 3.4 and Corollary 3.5, we can find a spin filling $X^{\prime}$ of the $X_{2}$-boundary component of $X$ in which $J_{2}$ bounds a characteristic disk $D$ with $D\cdot D\equiv 0$ mod 16. So $X^{\prime}$ is a spin filling of $S^{3}$ in which $C\cup D$ is a characteristic disk with self-intersection 0 mod 16 bounded by $J_{1}$. But $\operatorname{Arf}(J_{1})=1$, which contradicts Corollary 3.5. ∎
It is worth spelling out the following corollary to our above work, in the precise wording that we will eventually need in Section 5.
Corollary 3.7.
Let $K$ be a null-homologous knot in a closed 3-manifold $Y$. Suppose that $\operatorname{Arf}(K,Y)=1$. Then there does not exist any pair $(D,X)$ where $X$ is a spin 4-manifold with $\partial X=Y$ and $D$ is a properly embedded, characteristic disk in $X$ with $D\cdot D\equiv 0$ mod 16 such that $D=\partial K$.
3.3. Extension to arcs and additivity
This invariant can be easily extended to properly embedded arcs in 3-manifolds with spherical boundary that are null-homologous rel. boundary.
Definition 3.8.
Let $A$ be an arc properly embedded in a 3-manifold $Y$ such that $\partial Y\cong S^{2}$ and $A$ is null-homologous in $Y$ rel. boundary. Consider the union of $(A,Y)$ with $(\ell,B^{3})$ along their common boundary, where $\ell$ denotes the unknotted arc in $B^{3}$. We define $\operatorname{Arf}(A,Y)$ to be $\operatorname{Arf}(A^{\prime},Y^{\prime})$, where $A^{\prime}=A\cup\ell$ and $Y^{\prime}=Y\cup B^{3}$.
This is easily seen to be well-defined, as any two arcs that are null-homologous in their component 3-manifolds glue together to form a null-homologous knot in the union. We also note the following additivity property, which we will need later.
Lemma 3.9.
Let $K_{1}$ and $K_{2}$ be null-homologous knots in 3-manifolds $Y_{1}$ and $Y_{2}$, respectively. Choose 3-balls $B_{i}$ embedded in $Y_{i}$ such that $B_{i}$ intersects $K_{i}$ in a trivial tangle, and form the connected sum of pairs $(K,Y)=(K_{1}\#K_{2},Y_{1}\#Y_{2})$ along the boundary of $Y_{i}\setminus B_{i}$. Then:
$\operatorname{Arf}(K,Y)=\operatorname{Arf}(K_{1},Y_{1})+\operatorname{Arf}(K_{2},Y_{2})\mod 2.$
Proof.
Let $X_{1}$ and $X_{2}$ be spin cobordisms of $Y_{1}$ and $Y_{2}$ to $S^{3}$, with embedded Arf annuli for $K_{1}$ and $K_{2}$ terminating in knots $J_{1}$ and $J_{2}$, respectively. Then so long as the balls $B_{i}$ are sufficiently small, we may remove their trace through the cobordisms $X_{i}$ to obtain spin cobordisms $X_{i}^{\prime}$ between $Y_{i}\setminus B_{i}$ and $S^{3}\setminus B^{3}$, with embedded copies of $I\times I$ cobounded by the tangles of $K_{i}$ and $J_{i}$ together with the trace of the intersections of $K_{i}$ with $\partial B_{i}$. We then glue these two cobordisms with boundaries together to form a 4-manifold $X$ with boundary $Y\sqcup S^{3}$, with an annulus $C$ cobounded by $K$ in $Y$ and $J:=J_{1}\#J_{2}$ in $S^{3}$. It is then simple to check that $C$ is an Arf annulus, and so the result follows by the additivity of the classical Arf invariant. ∎
The above proof follows exactly the same for null-homologous arcs in 3-manifolds with spherical boundary.
4. Equivariantly slice 2-knots
In Section 2, we constructed for each periodic 2-knot a corresponding periodic Seifert 3-manifold $Y$ such that $\operatorname{Fix}(\tau)\cap Y=A$, with $A$ a single arc. Together with the definitions established in the previous section, we may define our invariant for periodic 2-knots.
Definition 4.1.
Let $S$ be a $d$-periodic 2-knot in $S^{4}$ such that $S$ bounds an equivariant Seifert 3-manifold $Y$ with fixed set $A$. Let $\overline{Y}$ be the quotient of $Y$ by $\tau$, and let $\overline{A}$ be the image of $A$ under the quotient map. The quantity $\alpha$ of $(S,Y,A)$ is then defined by:
$\alpha(S,Y,A):=\operatorname{Arf}(\overline{A},\overline{Y})\in\mathbb{Z}/2\mathbb{Z}$
We have not proven at this point that $\alpha$ is actually an invariant of $S$. The proof of this fact will follow from work in Section 5. However, we will be able to show in this section the following theorem, which states that periodic knots for which $\alpha$ vanishes for some Seifert 3-manifold are indeed slice.
Theorem 4.2.
Let $S$ be a $d$-periodic 2-knot, and let $Y$ be an equivariant Seifert 3-manifold with fixed arc $A$. If $\alpha(S,Y,A)=0$, then $S$ is equivariantly slice, regardless of the order $d$ of $\tau$.
The strategy to constructing the slice ball is as follows: First, we quotient $Y$ by $\tau$ to obtain some other 3-manifold $\overline{Y}$, as in the definition of $\alpha$. This realizes $Y$ as the cyclic cover of $\overline{Y}$ branched over an arc $\overline{A}$, where $A$ descends to $\overline{A}$ under the branched covering map. We then isotope $\overline{Y}$ into the interior of $\overline{B^{5}}$ and ambiently surger the pair $(\overline{Y},\overline{A})$ to $(\overline{B^{3}},\overline{\ell})$ where $\overline{\ell}$ is an unknotted arc. In order to ensure that these surgeries can be performed ambiently without introducing new intersections with $\operatorname{Branch}(\tau)$, we will require $\alpha$ to vanish. We then lift the isotopy and surgery curves and disks to $Y\subset B^{5}$, obtaining a sequence of suitable equivariant surgeries taking $(Y,A)$ to $(B^{3},\ell)$, since the latter is its own cyclic branched cover.
We now proceed with the proof of Theorem 4.2, which will occupy the next few subsections. The first step of this procedure is to take the quotient of $Y$ by $\tau$. Note that by construction, $Y\cap\operatorname{Fix}(\tau)=A$, where $A$ is a single arc. Thus the quotient $Z:=Y/(y\sim\tau(y))$ is a topological 3-manifold $\overline{Y}$. In particular, if $\pi$ is the quotient map $Y\to Z$, then the $n$-fold cyclic cover of $\overline{Y}$ branched along $B:=\pi(A)$ recovers $Y$.
The following lemma allows us to push $\overline{Y}$ into $\overline{B^{5}}$, and by extension $Y$ into $B^{5}$, to prepare it for surgery. We will not actually use the extra properties of equivariant Morse position and the handle structure of the fixed set in this section, but it will be important later in Theorem 5.3. Note that in order to talk about equivariant isotopy of $Y$, we must realize it as the image of some equivariant map $\psi\colon M\to B^{5}$, where $M$ is an abstract 3-manifold of the same homeomorphism type as $Y$. Note $M$ then comes equipped with an action $\rho$ of $\mathbb{Z}/d\mathbb{Z}$ given by the pullback of $\tau$ by $\psi$.
Lemma 4.3.
For a fixed periodic 2-knot $S$, let $\psi\colon M\to B^{5}$ be a proper embedding whose image is the Seifert 3-manifold $Y$ as constructed in Proposition 2.1. Then $\psi$ is equivariantly isotopic rel. boundary to an equivariant proper embedding $\psi^{\prime}$ such that there exists an equivariant Morse function $h$ on $B^{5}$ with respect to which $\psi^{\prime}(M)$ is in ordered equivariant Morse position and $h|_{A}$ has only one critical point, where $A$ denotes $\psi^{\prime}(M)\cap\operatorname{Fix}(\tau)$.
Proof.
Let $\overline{M}$ be the quotient of $M$ by $\tau$, and let $\overline{\psi}$ be the induced map to $\overline{B^{5}}$ with image $\overline{Y}$. Choose a relative Morse function $g$ on $\overline{M}$ (which is by definition identically 0 on $\partial M$) which is valued in $[0,1]$ achieves its maximum on an interior point of $\overline{\psi}^{-1}(\overline{A})$.
Fix a local equivariant Morse function $\overline{h}$ on a neighborhood $U$ of $\overline{Y}$ in $\overline{B^{5}}$ (including some of the interior of $B^{5}$, not just on the boundary) such that $h|_{\overline{\partial B^{5}}}\equiv 0$. We can then construct an isotopy $\overline{\Psi}\colon\overline{M}\times[0,1]\to h^{-1}([0,1])$ given by:
$\overline{\Psi}(x,t)=(\overline{\psi}(x),t\cdot g\circ\overline{\psi}(x))$
Then $\overline{\Psi}(x,0)=\overline{\psi}(x)$ in the bounding $\overline{S^{4}}$, and $\overline{\Psi}(x,1)$ is a proper embedding $\overline{Y}$ of $\overline{M}$ in $\overline{B^{5}}$ such that $\overline{\Psi}(\overline{A})$ has exactly one maximum. Here again we use Proposition 1.11 to rearrange $\overline{Y}$ via ambient isotopy to ordered equivariant Morse position without changing the single critical point of $\overline{A}$.
We may then take the $d$-fold branched cover of $\overline{B^{5}}$ and $\overline{Y^{\prime}}$ inside it to get $Y$ pushed into $B^{5}$ in ordered equivariant Morse position with respect to the lifted equivariant Morse function $h$ on $B^{5}$. Since $\overline{A}$ had only a single maximum with respect to $\overline{h}$, $A$ has only a single maximum with respect to $h$. As before, this isotopy lifts by Schwarz’s Isotopy Lifting Theorem [Sch80]. ∎
As the reader will see later in the following section, it is crucial that the above lemma allows us to preserve the single critical point of $\overline{A}$.
4.1. Spin structures and ambient surgery in dimension 5
We take a moment to describe a restraint on ambient surgeries for 3-manifolds in dimension 5. We adapt the language of Sunukjian [Sun15], although this obstruction was known to Kervaire and overcome in his original proof of sliceness of 2-knots. Note that for this discussion, we do not assume any equivariance.
Recall that a trivialization $\xi$ of a vector bundle $E$ over base $M^{n}$ with fiber $F^{k}$ is an identification of $E$ with $M\times F$. One can specify a trivialization directly with a map, or with a choice of $k$ linearly independent sections of $E$. We use the terms trivialization and framing interchangeably.
Recall that a spin structure on an $n$-manifold $M$ can be thought of as a trivialization of the tangent bundle $TM$ restricted to the 1-skeleton $M_{1}$ that extends over the 2-skeleton $M_{2}$. A discussion of the equivalence of this definition to the usual one involving lifts of principal $\text{SO}(n)$-bundles can be found in [Kir89, Ch. VI].
Now suppose that $Y$ is a 3-manifold, and let $\mathfrak{s}$ be a spin structure on $Y$. Let $\gamma$ be a curve in $Y$. We say that $\xi$-framed surgery along $\gamma$ is a spin surgery if the 4-manifold $X$ obtained by attaching a 4-dimensional 2-handle along $\gamma$ with framing $\xi$ to $Y\times\{1\}$ inside $Y\times[0,1]$ has a spin structure $\mathfrak{t}$ such that $\mathfrak{t}|_{Y\times\{0\}}=\mathfrak{s}$.
Finally, we recall the definition of ambient surgery. Let $Y$ be a 3-manifold embedded in $B^{5}$, and let $\gamma$ be a curve in $Y$ such that $\gamma$ bounds a disk $\Delta$ into $B^{5}\setminus Y$. Let $E$ be the total space of an embedded 2-disk sub-bundle of the normal bundle $\nu_{B^{5}}(\Delta)$ such that $\partial E\cap Y=\nu_{Y}(\gamma)$. Finally, let $\xi$ be a framing of $\nu_{Y}(\gamma)$ that extends to the unique trivialization of $E$ over $\Delta$. Then the manifold $Y^{\prime}$ embedded in $B^{5}$ given by $Y^{\prime}:=Y\setminus\nu_{Y}(\gamma)\cup(\Delta\times\partial D^{2})$ is the result of ambient surgery along $\gamma$ with framing $\xi$.
It is the purpose of the following lemma to determine when ambient surgery can be performed; that is, given $\gamma$ and $\Delta$, for which framings $\xi$ we can perform this operation.
Lemma 4.4.
Let $Y$ be a 3-manifold embedded in $B^{5}$, and let $\gamma$ be a curve in $Y$ with a choice of framing $\xi$ such that $\gamma$ bounds a disk $\Delta$ into $B^{5}\setminus Y$. Let $\mathfrak{s}$ be the unique spin structure on $B^{5}$. Let $Y^{\prime}$ be the result of $\xi$-framed surgery along $\gamma$. Then the surgery $Y$ to $Y^{\prime}$ can be realized as an ambient surgery in $B^{5}$ if and only if $\xi$-framed surgery along $\gamma$ is a spin surgery with respect to $\mathfrak{s}|_{Y}$.
Proof.
Since $\gamma$ is a curve in an orientable a 3-manifold, trivializations of its tubular neighborhood $\nu_{Y}(\gamma)$ are in bijection with $\pi_{1}\text{SO}(2)\cong\mathbb{Z}$ and carry an action of $\mathbb{Z}$ which increases or decreases the number of full counterclockwise twists. On the other hand, trivializations of $\nu_{B^{5}}(\gamma)$ are in bijection with $\pi_{1}\text{SO}(4)\cong\mathbb{Z}/2\mathbb{Z}$. Indeed, the inclusion map of $\nu_{Y}(\gamma)$ into $\nu_{B^{5}}(\gamma)$ induces the mod 2 map from $\mathbb{Z}$ to $\mathbb{Z}/2\mathbb{Z}$. Thus, if $\xi$ and $\xi^{\prime}$ are two different trivializations of $\nu_{Y}(\gamma)$ that differ by an even number of twists, then when included into $\nu_{B^{5}}(\gamma)$, they become homotopic.
Since $\Delta$ is contractible, every 2-disk bundle over it is canonically trivial; denote this trivialization $\zeta$. Thus given a 2-disk bundle over $\Delta$, we expect that half of the possible framings on $\gamma$ can be extended over the bundle, depending on their parity relative to $\zeta$.
Now $\Delta$ is embedded in $B^{5}$, and so the unique trivialization of its normal bundle $\nu_{B^{5}}(D)$ is exactly the unique spin structure $\mathfrak{s}$ on $B^{5}$ restricted to $\Delta$. This then further restricts to a trivialization $\zeta$ on any 2-disk sub-bundle of $\nu_{B^{5}}(\Delta)$, and in particular on $\nu_{B^{5}}(\partial\Delta)$, which is just $\nu_{B^{5}}(\gamma)$. Thus the framings $\xi$ on $\nu_{Y}(\gamma)$ for which there exist embedded $D^{2}$ sub-bundles of $\nu_{B^{5}}(\Delta)$ over which they can be extended are precisely those that agree with $\zeta$ restricted to $\nu_{B^{5}}(\gamma)$. Since $\zeta$ was induced by the restriction of the ambient spin structure on $B^{5}$, we see that surgery along $\gamma$ with framing $\xi$ must be spin surgery. ∎
4.2. Equivariant surgery preserving the fixed set
Here we address another restraint on the allowable surgeries that does not occur in non-equivariant surgery. Suppose we wished to surger along a curve $\gamma$ in $\overline{Y}$ that did not bound any disk disjoint from the image of $\operatorname{Branch}(\tau)$ in the quotient $\overline{B^{5}}$. This will certainly be the case if $\gamma$ lifts to a single curve, but even if it lifts to two curves, if these curves are not null-homotopic in $B^{5}\setminus(Y\cup\operatorname{Fix}(\tau))$, then when we try to surger $Y$ along them, we will introduce new connected components of the fixed set of our surgered 3-manifold. This is illustrated in Figure 8.
Figure 8. New components of the fixed set introduced by errant handle attachment. On the left, we see $\gamma$ (horizontal) linking $\operatorname{Fix}(\tau)$ (vertical) inside $Y$. On the right, we see a different 3-dimensional projection of the same intersection, in which $\gamma$ appears as a pair of points and $\operatorname{Fix}(\tau)$ is a plane. The arc connecting these two points and intersecting $\operatorname{Fix}(\tau)$ is one slice of the disk $\Delta$ that we want to surger along, with its normal neighborhood shown in blue. However, as we see, this neighborhood intersects $\operatorname{Fix}(\tau)$, and so surgery along $\gamma$ will introduce a new component of the fixed set of $Y$ given by the intersection of the boundary of the tubular neighborhood of $\Delta$ with $\operatorname{Fix}(\tau)$.
It is a theorem of Birman-Hilden [BH75] together with Morgan’s work on the Smith Conjecture [Mor84] that if the cyclic branched cover of $S^{3}$ over any link is itself $S^{3}$, then the link is the unknot. Thus our only hope for surgering $Y$ to $B^{3}$ is to make sure our surgeries on $\overline{Y}$ do not introduce new components of $\overline{A}$. Therefore, in order to check if a curve $\gamma$ in $\overline{Y}$ can be surgered along, we need a way to check that it bounds a disk that does not intersect either $\overline{Y}$ or $\operatorname{Branch}(\tau)$. The following proposition will help us do this.
Proposition 4.5.
[…]
Proof.
We make a rising waters argument; compare with [GS00], Prop. 6.2.1 for this type of argument for surfaces in 4-manifolds. Consider the equivariant radial Morse function $h$ on $B^{5}$ from Lemma 4.3. Note that both $Y$ and $\operatorname{Fix}(\tau)$ are properly embedded and boundary parallel, and their intersection is a single properly embedded tangle. Furthermore, $\operatorname{Fix}(\tau)$ consists of exactly one index 0 critical point with respect to $h$, and so does the intersection $\cap\operatorname{Fix}(\tau). YY$ itself has a unique index 0 critical point, but may have other critical points as well.
The minima of $Y$ and $\operatorname{Fix}(\tau)$ with respect to this Morse function both contribute a 1-handle each to the complement. The minimum point of their intersection then contributes a 2-handle, the attaching sphere of which winds around the commutator of the meridians of both $Y$ and $\operatorname{Fix}(\tau)$ near the intersection point. Thus the fundamental group of the sublevel set of $h$ just above the first intersection point is $\mathbb{Z}^{2}$. However, all subsequent critical points of $Y\cup\operatorname{Fix}(\tau)$ are in $Y$ and are of index 1 or 2, which contribute 2- or 3-handles to the complement. Of these, only the 2-handles can contribute to the fundamental group relations of $B^{5}\setminus(Y\cup\operatorname{Fix}(\tau))$, but these only introduce extra commutation relations between meridians that already commute. ∎
We now pass to the quotient under $\tau$, in which the meridians of $Y$ and $\operatorname{Fix}(\tau)$ are mapped to meridians of $\overline{Y}$ and $\operatorname{Branch}(\tau)$, respectively, and no new relations are added. Thus any curve that is null-homologous in $\overline{B^{5}}\setminus(Y\cup\operatorname{Branch}(\tau))$ is also null-homotopic in $\overline{B^{5}}\setminus(\overline{Y}\cup\operatorname{Branch}(\tau))$. Hence we conclude that if $\gamma$ is a curve in $\overline{Y}$ that is null-homologous in $\overline{B^{5}}\setminus\operatorname{Branch}(\tau)$, then it bounds a disk in $\overline{B^{5}}\setminus(\overline{Y}\cup\operatorname{Branch}(\tau))$ and we may therefore surger along it.
4.3. Finding surgery curves
Now to ambiently surger $(\overline{Y},\overline{A})$ to $(\overline{B^{3}},\overline{\ell})$, we proceed in two steps. The first is to ignore $\overline{A}$ and simply find surgery curves that realize a spin cobordism between $\overline{Y}$ and $\overline{B^{3}}$. We can find such curves by Kaplan’s theorem [Kap79]. Once we identify the curves we wish to surger $\overline{Y}$ along, we isotope them inside $\overline{Y}$ so that they are null-homologous in $B^{5}\setminus(Y\cup\operatorname{Fix}(\tau))$, allowing them to pass through $\overline{A}$ if necessary. Such an isotopy inside $\overline{Y}$ can always be found as $\overline{A}$ is null-homologous in $\overline{Y}$ rel. boundary, and so if a surgery curve $\gamma$ in $\overline{Y}$ does not bound a disk in $\overline{B^{5}}\setminus\overline{Y}\cup\operatorname{Fix}(\tau)$, we may use a surface in $\overline{Y}$ bounded by $\overline{A}$ to isotope it along a path to $\overline{A}$ and push it past $\overline{A}$ to remove the intersection. The result of the surgery on these curves will be to turn $\overline{Y}$ into $\overline{B^{3}}$, realizing $\overline{A}$ as an arc in a 3-ball.
It remains then to unknot $\overline{A}$ via surgery preserving $B^{3}$. We do this diagrammatically with explicit surgery curves that are null-homologous in the complement of $\overline{A}$. To do this, choose an arc in $S$ connecting the two boundary points of $\overline{A}$, and let $B^{\prime}$ be the knot obtained as the union of this arc with $\overline{A}$, pushed into the interior of $B^{3}$. Recall by Lemma 3.5 that if $\operatorname{Arf}(B^{\prime})=0$, then $B^{\prime}$ is band pass equivalent to the unknot, and that these band passes can be achieved by 0-framed (thus spin) surgery on null-homologous curves, as shown in Figure 9. If $\operatorname{Arf}(B^{\prime})=1$, then we can use Lemma 3.2 to surger $\overline{A}$ to the trefoil tangle if we wish.
[…]
Figure 9. A band pass achieved by 0-framed 2-handles.
4.4. Lifting the surgeries
Let $\{\gamma_{i},\Delta_{i}\}$ be the collection of framed curves $\gamma$ in $\overline{Y}$ that we wish to surger along, and let $\Delta_{i}$ be the framed disks that they bound, as found in the previous subsection. Then since all the $\gamma_{i}$ are null-homotopic in the complement of the branch set, they lift to $n$-tuples of null-homotopic curves in the complement of $\operatorname{Fix}(\tau)$. The disks $\Delta_{i}$ lift as well, and are bounded by the lifts of the $\gamma_{i}$. Finally, since surgery on the $\gamma_{i}$ in $\overline{Y}$ gave $\overline{B^{3}}$ with fixed set consisting of the trivial arc $\overline{\ell}$, we conclude that the lifts of these surgeries equivariantly surger $Y$ to $B^{3}$. This concludes the proof of Theorem 4.2. ∎
We may also leverage the construction in this proof to obtain a concordance statement about 2-knots for whom we can construct an equivariant Seifert 3-manifold $Y$ with fixed arc $A$ such that $\operatorname{Arf}(\overline{A},\overline{Y})=1$.
Proposition 4.6.
Let $S$ be a $d$-periodic 2-knot with equivariant Seifert 3-manifold $Y$ with fixed arc $A$. If $\alpha(S,Y,A)=1$, then $S$ is equivariantly concordant to the $d$-twist spun trefoil. In particular, if $d$ is odd, then $S$ bounds an equivariant homology ball.
Proof.
By the construction above and the fact that every knot is band pass equivalent to either the unknot or the trefoil, we see that for any triple $S,Y,A$, with $\operatorname{Arf}(\overline{A},\overline{Y})=1$, we may ambiently surger $\overline{Y}$ to $B^{3}$ and $\overline{A}$ to the trefoil without introducing new components of the fixed set. Zeeman [Zee65] showed that the $d$-twist spin of any classical knot $K$ bounds a punctured copy of the $d$-fold cyclic cover of $S^{3}$ branched over $K$. Thus the quotient of such a Seifert solid by the rotation action permuting the twists retrieves $B^{3}$ branched along the tangle $K$. Hence, the surgeries then lift to a sequence of surgeries from $Y$ to the $d$-fold branched cover of $B^{3}$ over the trefoil, which is bounded by the $d$-twist spun trefoil.
For the latter statement, we modify this construction slightly. For integers $p,q,$ and $d$ are pairwise coprime, the Brieskorn homology sphere $\Sigma(p,q,d)$ is the $d$-fold branched cover of $S^{3}$ over the torus knot $T(p,q)$. The knot $T(p,q)$ has Arf invariant 1 if and only if $p$ is even and $q\equiv\pm 3$ mod 8; this follows from work of Levine [Lev66]. Thus if $\tau$ has order $d$ and $\operatorname{Arf}(\overline{A},\overline{Y})=1$, then we may choose an appropriate $p$ and $q$ such that $(\overline{A},\overline{Y})$ can be surgered to $(\overline{T},\overline{B^{3}})$, where $\overline{T}$ denotes the $T(p,q)$-tangle in the quotient $\overline{B^{5}}$. After this surgery, we may take the $d$-fold branched cover to see that $S$ bounds a copy of $\Sigma(p,q,d)$. ∎
At this point, we have shown that there are at most two equivariant concordance classes of $d$-periodic 2-knots: one containing the unknot, and the other containing the $d$-twist spun trefoil. However, we have not yet shown that the quantity $\alpha$ defined in Definition 4.1 is a concordance invariant of periodic 2-knots, and so we cannot say yet that these two concordance classes are not the same. In the following section, we will prove these remaining statements.
5. An obstruction to equivariant sliceness
The goal of this section is to make use of Lemma 3.5 to obstruct the existence of equivariant slice balls for 2-knots for which there exists a quotient Seifert 3-manifold with branch sets with Arf invariant 1. The first step to this is, in fact, showing the existence of $\tau$-equivariant Seifert 4-manifolds for simply periodic 3-manifolds in $S^{5}$.
5.1. Equivariant Seifert 4-manifolds
Proposition 5.1.
Let $Y$ be a simply $d$-periodic 3-manifold equivariantly embedded in $S^{5}$ with respect to the standard order $d$ rotation action $\tau$ fixing an embedded copy of $S^{3}$ such that $K:=Y\cap\operatorname{Fix}(\tau)$ is a copy of $S^{1}$. Suppose there exists an equivariant Morse function $h$ on $S^{5}$ such that $Y$ is in ordered equivariant Morse position with respect to $h$, and $h|_{K}$ is Morse and has exactly two critical points and is fully vertical with respect to $h$ away from these critical points. Then $Y$ bounds an equivariantly embedded simply $d$-periodic 4-manifold $X\subset S^{5}$ such that $X\cap\operatorname{Fix}(\tau)$ is a characteristic disk $D$ properly embedded in $X$ with $D\cdot D=0$.
Proof.
We follow roughly the same strategy as in the proof of Proposition 2.1. That is, we build $X$ handle by handle according to an equivariant handle decomposition of $Y$ coming from $h$. We do this, as before, by finding equivariant fillings for $t$-slices of $Y$, which themselves are equivariant surfaces. To better visualize these surfaces, we parametrize a neighborhood of $Y$ so that $\operatorname{Fix}(\tau)$ appears locally as the $(z,w,t)$-subspace, and so that $h$ is just projection to the $t$ coordinate. Note that the lowest minimum (maximum) of $Y$ is necessarily also the minimum (resp. maximum) of $K$, and so every $t$-slice has at least one fixed point in it. The assumption that $K$ is fully vertical with respect to $h$ away from its critical points implies that no $t$-slice has more than two fixed points in it. Thus in each $t$-slice apart from those containing the very lowest minimum and the very highest maximum, $Y$ appears as a surface in 4-space, equivariant with respect to rotation about the $(z,w)$-axis, with two fixed points. Our goal is therefore to construct smoothly varying 3-dimensional fillings $X_{t}$ of these surfaces $Y_{t}$ such that each of these slices intersects $\operatorname{Fix}(\tau)$ in a single smoothly varying arc. These arcs stacked on top of each other will then form a disk, which will be $\operatorname{Fix}(\tau)\cap X$ as desired. As before, we provide a schematic showing the handle attachments at different stages.
[…]
Figure 10. A schematic showing the attachment of handles of $X$ with respect to the Morse function $h$.
$\boldsymbol{f=0.}$ Since $Y$ is in equivariant Morse position, we can assume that all of its 0-handles occur in $d$-tuples, except for one, which is equivariant with respect to $\tau$. We may begin constructing $X$ by filling in these 0-handles with 4-dimensional index 0 right half-handles. This completes $X$ up to $t=\frac{1}{2}$.
$\boldsymbol{f=\frac{1}{2}.}$ At this point, $X_{\nicefrac{{1}}{{2}}}$ is an equivariant collection of 3-balls in ambient 4-space. As in the previous case, we want to fill the 1-handles of $Y$ with 4-dimensional right half-handles. However, the trajectories from the index 1 critical points of $Y$ may link slices of the 0-handles of $Y$. As before, we consider these trajectories intersecting the 0-handles of $Y$ in $h^{-1}(\frac{1}{2},\frac{1}{2}+\epsilon)$. To avoid these intersections, we preemptively attach interior 1-handles to $X$ at $t=\frac{1}{2}-\epsilon$ along the trajectories to create 3-dimensional tubes for these trajectories to pass through.
$\boldsymbol{f=1.}$ After these interior 1-handles are attached, we may proceed to attach the 3-dimensional 1-handles and their 4-dimensional fillings.
$\boldsymbol{f=\frac{3}{2}.}$ Following the same general strategy, we now examine the descending manifolds of the 3-dimensional index 2 critical points at $t=2$. Because $Y\cap\operatorname{Fix}(\tau)$ is precisely $K$, we know none of these descending manifolds intersect $\operatorname{Fix}(\tau)$, as all such intersections are already accounted for by our assumption that $K$ has precisely two critical points with respect to the ambient Morse function. Thus we may safely attach right half-handles to $X$ for each 2-handle of $Y$ without fear of breaking equivariance or introducing new components of the fixed set. However, the index 2 descending manifolds may intersect interior 1-handles of $X$ that we have already attached. As before, suppose that all of these intersections occur within the preimage $h^{-1}(\frac{3}{2},\frac{3}{2}+\epsilon)$.
To avoid these intersections, note that the generic intersection between a 2-handle core and a 4-dimensional 1-handle in $B^{5}$ is a disk in the 2-handle core. Thus we can use this disk as a guide to preemptively attach 4-dimensional interior 2-handles to compress the 4-dimensional 1-handles before attaching the 3-dimensional 2-handle. This is completely analogous to the preemptive 1-handles we attached to make way for the 1-handles of $Y$. Note further that because the disks we use as guides for attaching our 4-dimensional 2-handles are already embedded in $B^{5}$ as the cores of existing 3-dimensional 2-handles, they have trivial normal bundles. Thus from the perspective of $X$, the compressions are achieved by attaching interior 2-handles, with framings given by a choice of trivial disk sub-bundle of the normal bundle. Furthermore, since none of the 2-handles of $Y$ intersect $\operatorname{Fix}(\tau)$, we may arrange so that the parallel 2-handles of $X$ do not intersect $\operatorname{Fix}(\tau)$ either.
$\boldsymbol{f=2.}$ Once we have performed these compressions at $t=\frac{3}{2}-\epsilon$, we may safely proceed upwards in $t$ and attach the 3-dimensional 2-handles of $Y$ and their 4-dimensional half-handle fillings at $t=2$.
$\boldsymbol{f=\frac{5}{2}.}$ At this point, for $t\in(2+\epsilon,\frac{5}{2}-\epsilon)$, $Y_{t}$ appears as an equivariant unlink of unknotted spheres with a pair of fixed points on a single component, and $X_{t}$ appears as a connected equivariant 3-manifold with boundary $Y_{t}$ and fixed set a single, “straight” line (with respect to our parameterizing coordinates). Now every component of $Y_{\nicefrac{{5}}{{2}}-\epsilon}$ bounds a 3-ball embedded in $B^{5}_{\geq 2+\epsilon}$ given by the descending manifolds of each maximum of $h$ on $Y$. As in the 3-manifold case, we may isotope these 3-ball caps straight down into the $t=\frac{5}{2}-\epsilon$-slice, using their boundary unknotted 2-spheres as guides near the boundary. We may also assume that these 3-balls are all equivariant, meaning that one central 3-ball intersects $\operatorname{Fix}(\tau)$ in a straight arc connecting the fixed points of $Y_{2+\epsilon}$ and the rest are permuted by $\tau$.
As in the lower-dimensional case, the boundaries of tubular neighborhoods of each of these 3-balls then form splitting spheres for $Y_{\nicefrac{{5}}{{2}}-\epsilon}$. These splitting spheres intersect $X_{\nicefrac{{5}}{{2}}-\epsilon}$ in a collection of 3-balls, which we may then use as equivariant compressing 3-balls for $X_{\nicefrac{{5}}{{2}}-\epsilon}$. We can then equivariantly compress $X$ along these 3-balls by attaching interior 3-handles to $X_{\nicefrac{{5}}{{2}}}$ and filling in closed components.
$\boldsymbol{f=3.}$ In fact, because the splitting spheres we found were the boundaries of small neighborhoods of balls bounded by components of $Y_{\nicefrac{{5}}{{2}}}$, each component of $X_{\nicefrac{{5}}{{2}}}$ with boundary is itself a 3-ball. We may subsequently fill the remaining closed components with equivariant 4-manifold fillings. Finally, we trace the isotopy of the 3-ball caps up to the maxima of $Y$ to find equivariant 4-balls that fill these maxima. This completes the construction of $X$ and the proof of Proposition 5.1. ∎
Remark 5.2.
The approach used in Propositions 2.1 and 5.1 is indicative of a more general strategy for building embedded Seifert $n$-manifolds inside $B^{n+1}$ explicitly handle by handle. That is, for every $k$-handle attachment of the boundary, attach interior $k$-handles to the Seifert filling to make way for the descending manifolds of the corresponding critical points. This approach was necessary for this paper in order to control the intersection of the the Seifert manifolds with the ambient fixed set.
5.2. Proof of obstruction
Recall that for a periodic 2-knot $S$ with equivariant Seifert 3-manifold $Y$ such that $Y\cap\operatorname{Fix}(\tau)$ is an arc $A$, the quantity $\alpha(S,Y,A)$ was defined to be:
$\alpha(S,Y,A)=\operatorname{Arf}(\overline{A},\overline{Y})$
where $\overline{A}$ and $\overline{Y}$ denote the image of these manifolds after taking the quotient by $\tau$.
Theorem 5.3.
Let $S$ be a periodic 2-knot and suppose $Y_{1}$ and $Y_{2}$ are two quotients of equivariant Seifert 3-manifolds with fixed arcs $A_{1}$ and $A_{2}$, respectively. Then $\alpha(S,Y_{1},A_{1})=\alpha(S,Y_{2},A_{2})$.
Proof.
Begin by embedding $S$ in $S^{4}$, thought of as an equator of $S^{5}$, so that the $S^{4}$ bounds two equivariant 5-balls whose union is all of $S^{5}$. Note that $\tau$ can then be extended to $S^{5}$ by gluing the actions on each $B^{5}$ on either side of the equator together, so that the action on $S^{5}$ fixes a 3-sphere. Parametrize a neighborhood of $S$ in $S^{5}$ with our usual conventions, though we now allow for $t$ to take negative values in order to describe points in both 5-balls.
Embed $Y_{1}$ in the $t\geq 0$ copy of $B^{5}$ such that $A_{1}$ has exactly one maximum with respect to $t$. In the other copy of $B^{5}$, embed $Y_{2}$ such that $A_{2}$ has exactly one minimum with respect to $t$. Note that these embeddings can be found as a result of Lemma 4.3.
The union $Y$ of $Y_{1}$ and $Y_{2}$ along $S$ is then a simply periodic closed 3-manifold, and its fixed set under $\tau$ is the union $K$ of $A_{1}$ with $A_{2}$. Note that with respect to projection onto the $t$-axis, $K$ has exactly two critical points. The pair $Y,K$ then satisfies the conditions of Theorem 5.1. Therefore, $Y$ bounds an equivariant, spin Seifert 4-manifold $X$ with fixed set a characteristic, 0-framed disk $D$ with boundary $K$.
We now take the quotient of this whole construction by $\tau$. Note that $\overline{D}$ remains characteristic and 0-framed in $\overline{X}$. Thus by Corollary 3.7, $\operatorname{Arf}(\overline{K},\overline{Y})$ must vanish. Hence by the additivity of the Arf invariant proved in Lemma 3.9, we conclude that $\alpha(S,Y_{1},A_{1})=\alpha(S,Y_{2},A_{2})$. ∎
Having concluded this, we see that $\alpha$ is independent of the choice of equivariant Seifert solid $Y$, and so it is an invariant of periodic 2-knots. Thus by Theorem 4.2 and Proposition 4.6, we have shown that the invariant $\alpha$ completely classifies $d$-periodic 2-knots up to equivariant concordance. As an immediate corollary, we obtain the following.
Theorem 5.4.
For all $d\geq 2$, $\mathcal{C}^{2}_{d}\cong\mathbb{Z}/2\mathbb{Z}$, generated by the $d$-twist spun trefoil.
Proof.
Consider the map $\varphi$ from $\mathcal{C}^{2}_{d}$ to $\mathbb{Z}/2\mathbb{Z}$ given by $S]\mapsto\alpha(S)$. By Theorem 5.3, $\alpha$ is well-defined on equivariant concordance classes, and so $\varphi$ is a well-defined map. The combination of Theorem 4.2 and Proposition 4.6 show that $\varphi$ is injective, as we can realize an equivariant concordance between any $d$-periodic 2-knot and either the unknot or the $d$-twist spun trefoil. We have also seen in the proof of Proposition 4.6 that $\alpha$, and thus $\varphi$, takes the value 1 on the $d$-twist spun trefoil for all $d$.
Finally, to show that $\varphi$ is indeed a homomorphism, we appeal to the additivity of the Arf invariant. Let $S$ and $S^{\prime}$ be two periodic 2-knots with equivariant Seifert 3-manifolds $Y$ and $Y^{\prime}$. Then their equivariant connected sum $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$ bounds an equivariant Seifert 3-manifold $Y\natural_{\mathbb{Z}/d\mathbb{Z}}Y$ whose fixed set is the connected sum of the fixed sets of $Y$ and $Y^{\prime}$ as arcs. By Lemma 3.9, the Arf invariant of the connected sum of the fixed sets is the sum of the individual Arf invariants mod 2. Thus $\alpha(S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime})=\alpha(S)+\alpha(S^{\prime})$. This proves that $\varphi$ is an isomorphism, and thus that $\mathcal{C}^{2}_{d}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$. ∎